# On Entropy Control in LLM-RL Algorithms

### v1

@@ 4.2 Adaptive clamped entropy control
where $\beta$ is the coefficient learning rate, and $\tilde{{\mathcal{H}}}_{\rm low},\tilde{{\mathcal{H}}}_{\rm high}$ are respectively the lower and upper limit of the (clamped) entropy. The algorithm will try to confine $\tilde{{\mathcal{H}}}$ within $\tilde{{\mathcal{H}}}_{\rm low},\tilde{{\mathcal{H}}}_{\rm high}]$ by increasing/decreasing $\lambda$ when $\tilde{{\mathcal{H}}}(\pi_{\theta})$ is lower/higher than the limits. The intuition is that when entropy is high, the coefficient should be tuned down to reduce the entropy induced bias and shift weights to greedy reward maximization, which in turn consumes entropy. While when entropy level is too low, the coefficient can be tuned up to leverage the benefits of entropy regularization. For better training stability, the entropy coefficient is also boxed in the range $\lambda_{\rm low},\lambda_{\rm high}]$ so that large fluctuations of entropy do not lead to coefficient over-shoot. Empirically, we find that such an update rule can effectively confine the policy entropy, which can further improve reasoning efficiency by avoiding entropy and response length explosion.
@@ 4.3 Algorithm
where ${\mathcal{L}}_{\rm PO}(\theta)$ is a policy optimization objective of which the gradient is an ascent direction of $V^{\pi_{\theta}}$. Candidates of ${\mathcal{L}}_{\rm PO}$ include the objectives introduced in Section 2.1, e.g., GRPO objective is used in our tests. At each global step, we set the sampling policy $\pi_{b}$ according to the choice of policy optimization objective ${\mathcal{L}}_{\rm PO}$. For example, in PPO-type algorithms, $\pi_{b}$ is set as $\pi_{\rm old}$ which is the policy from last global step. Given $\pi_{b}$, a batch of queries $s_{0}\sim{\mathcal{D}}$ are sampled, and for each query, $\pi_{b}$ rolls out a batch of trajectories up to the maximum time step. With the batched samples, we can then optimize for the batch surrogate of ${\mathcal{L}}_{\rm AEnt}(\theta;\lambda)$ for several mini-epochs. At the end of each global step, the entropy coefficient is adjusted according to scheme 4.1. The whole process is summarized in Algorithm 1.
@@ 5.1 Training details
Evaluation. For performance comparison, we evaluate models on the AIME 2024, MATH-Hard test split, MATH-500, AMC23, MinervaMath (Lewkowycz et al., 2022) and OlympiadBench (He et al., 2024). We estimate the test score by averaging 4 tries per query on all benchmarks. The test-time generation temperature is $0.6$, top-p is $0.95$ and top-k is $20$.
Hyper-parameter settings. The tests are based on the verl framework (Sheng et al., 2025) [^1] The code will be released as soon as an internal mandatory pre-open-source review is completed.. When training Qwen2.5-math-1.5b base model on the MATH dataset, we use AdamW optimizer with a learning rate of $2\times 10^{-6}$. We set the max response length as $3072$. We use a batch size of $512$, and for each query we roll out $16$ responses with default sampling parameters (top-p and temperature set as $1$). For AEnt, we use the GRPO loss as ${\mathcal{L}}_{\rm PO}$. We use a clamping percentage $p=0.33$, and set $\tilde{{\mathcal{H}}}_{\rm low}\!=\!0.13$ and $\tilde{{\mathcal{H}}}_{\rm high}\!=\!0.24$. We use an initial entropy coefficient of $0.002$, and start updating the coefficient from the third epoch with a learning rate of $0.002$. We clip the coefficient in between $0.0006$ and $0.009$. For EntReg method, we use the traditional entropy bonus with a fixed entropy coefficient of $0.002$. When training DeepSeek-R1-distilled-Qwen model on the OpenR1-math dataset, we use a learning rate of $1\!\times\!10^{-6}$, a max response length of $7168$, a batch size of $256$ and for each query we roll out $8$ responses. We use $p=0.25$, $\tilde{{\mathcal{H}}}_{\rm low}\!=\!0.35$ and $\tilde{{\mathcal{H}}}_{\rm high}\!=\!0.62$, an initial coefficient of $3\!\times\!10^{-4}$, and start updating the coefficient from the second epoch with a learning rate of $10^{-4}$. We clip the coefficient in between $4\times 10^{-5}$ and $0.001$.
@@ 5.2 Performance analysis
An observation on the test score and the entropy trend. An interesting observation from Figures 4(a) is that after around $175$ steps (collapse time), the policy entropy of GRPO largely depletes and the entropy of EntReg starts to drastically fluctuate, while AEnt’s policy entropy is kept stable. Then one can observe from Figure 3(a) that the test score of GRPO and EntReg saturates around the same step, while the score of AEnt continues to improve and surpasses the baselines past the collapse time. This observation is consistent with our intuition and theoretical analysis: after GRPO’s entropy collapse, its policy becomes concentrated on few paths and no new information can be gained in the sampling process, ultimately leading to the stagnancy of the learning process. This is predicted by Proposition (I) that the policy will become stationary after entropy depletion.
@@ 5.3 Ablation studies
In this section, we conduct ablation studies on our algorithm.
Adaptive coefficient stabilizes training. In Figure 5, we compare the performance of adaptive coefficient vs constant coefficient. The test performance is similar for the two methods in this particular experiment. However, adaptive coefficient leads to a significant advantage on reasoning efficiency by delivering more compact responses while not sacrificing accuracy. In the third plot of Figure 6, constant coefficient fails to stabilize policy entropy in the mid of training, which results in the entropy blow up. We observe a positive correlation between entropy and response length in this case, where a exploding entropy leads to repeated reasoning patterns that do not increase the test scores. On the other hand, the adaptive coefficient successfully prevents the entropy and response length from blowing up.
Analysis of the entropy clamping percentage $p$. In Figure 6, we compare the algorithmic performance under different choice of clamping percentage $p$. It can be observed that a too large or too small $p$ can result in non-optimal gains. Intuitively, the percentage $p$ decides the size of the clamped space ${\mathcal{A}}(s)$, where a larger $p$ leads to more aggressive clamping and less tokens taken into account during entropy calculation. This would smooth the LLM policy on a more compact space, reducing the bias induced by entropy maximization while running the risk to leave out valuable tokens. In this sense, it is reasonable to try to maximize $p$ until the performance drops, which is also suggested by our reported results. Despite the fact the AEnt’s performance is affected by the choice of $p$, its advantage over the baselines is somewhat robust to the choice. It can be observed from Figure 6 that AEnt outperforms the baselines with different choices of $p$.
6 Conclusion and future work
@@ Appendix A Appendix
A.1 Preliminary lemmas
@@
A.2 Proof omitted in Section 3
A.2.1 Proof of Proposition 1
@@
Notice that the policy is an auto-regressive LLM and the transition is deterministic with $\mathbb{P}(s_{t+1}=(s_{t},a_{t})|s_{t},a_{t})=1$, we will have $\mathbb{P}_{t}^{\pi_{\theta}}(s_{h}^{*})=0$ for any $t\neq h$. Using this fact in equation A.26
@@
$\|\nabla V^{\pi_{\theta}}({\mathcal{D}})\|$ $\geq\frac{1}{\sqrt{H}|{\mathcal{D}}|}C^{\pi_{\theta}}(s_{0})\Big{(}V^{\pi^{*}}(s_{0})-V^{\pi_{\theta}}(s_{0})\Big{)}$ (A.31)
where $C^{\pi_{\theta}}(s_{0})=\max_{(a_{0},\dots,a_{H-1})\in{\mathcal{A}}_{H}^{*}(s_{0})}\Pi_{t=0}^{H-1}\pi_{\theta}(a_{t}|s_{t})$ with ${\mathcal{A}}_{H}^{*}(s_{0})=\{(a_{0},a_{1},\dots,a_{H-1})\in{\mathcal{A}}^{H}~|~\exists\pi^{*}\in\arg\max_{\pi}V^{\pi}({\mathcal{D}}),\Pi_{t=0}^{H-1}\pi^{*}(a_{t}|s_{t})>0\}$. ∎
A.3 Proof of Proposition 2
@@
$V_{\lambda}^{\pi_{\lambda}^{*}}(s_{0})-V_{\lambda}^{\pi_{\theta}}(s_{0})\leq\frac{1}{2\lambda}\frac{|{\mathcal{D}}|^{2}}{C_{\lambda}^{\pi_{\theta}}(s_{0})}\|\nabla V_{\lambda}^{\pi_{\theta}}({\mathcal{D}})\|^{2}$ (A.39)
@@
Define ${\mathcal{S}}(s_{0})\subseteq{\mathcal{S}}$ as the set of sequences starting from query $s_{0}$, i.e., ${\mathcal{S}}(s_{0})=\{s_{0},a_{0},\dots,a_{H-1}|\text{$a_{t}\in{\mathcal{A}}$ for $t\in[0,H-1]$}\}$. Then we have
@@
where the second and the third inequalities follow from Cauchy-Schwartz inequality. and the last inequality follows from equation A.3. The constant $C_{d}=\frac{1}{\sqrt{|{\mathcal{S}}(s_{0})|}}=\big{(}|{\mathcal{A}}|^{H}\big{)}^{-0.5}$.
@@
$\geq C_{d}^{2}\sum_{s\in{\mathcal{S}}(s_{0})}\sum_{t=0}^{H-1}(\mathbb{P}_{t}^{\pi_{\theta}}(s))^{2}(\min_{a}\pi_{\theta}(a|s))^{2}\Big{\|}Q_{t,\lambda}^{\pi_{\theta}}(s,\cdot)-\lambda\theta_{s,\cdot}-\frac{\sum_{a}Q_{t,\lambda}^{\pi_{\theta}}(s,a)-\lambda\theta_{s,a}}{|{\mathcal{A}}|}\Big{\|}_{\infty}^{2}$
$=C_{d}^{2}\lambda^{2}\sum_{s\in{\mathcal{S}}(s_{0})}\sum_{t=0}^{H-1}\mathbb{P}_{t}^{\pi_{\theta}}(s)(\min_{a}\pi_{\theta}(a|s))^{2}\frac{\mathbb{P}_{t}^{\pi_{\theta}}(s)}{\mathbb{P}_{t}^{\pi_{\lambda}^{*}}(s|s_{0})}\mathbb{P}_{t}^{\pi_{\lambda}^{*}}(s|s_{0})\Big{\|}Q_{t,\lambda}^{\pi_{\theta}}(s,\cdot)/\lambda-\theta_{s,\cdot}$
@@
$C_{\lambda}^{\pi_{\theta}}(s_{0})$ $=C_{d}^{2}\min_{t,s\in{\mathcal{S}}(s_{0})}\mathbb{P}_{t}^{\pi_{\theta}}(s)(\min_{s,a}\pi_{\theta}(a|s))^{2}\min_{t,s\in{\mathcal{S}}(s_{0})}\frac{\mathbb{P}_{t}^{\pi_{\theta}}(s)}{\mathbb{P}_{t}^{\pi_{\lambda}^{*}}(s|s_{0})}.$
@@
$V_{\lambda}^{\pi_{\lambda}^{*}}(s_{0})-V_{\lambda}^{\pi_{\theta}}(s_{0})\leq\frac{1}{2\lambda}\frac{|{\mathcal{D}}|^{2}}{C_{\lambda}^{\pi_{\theta}}(s_{0})}\|\nabla V_{\lambda}^{\pi_{\theta}}({\mathcal{D}})\|^{2}$ (A.49)
which completes the proof.

### v2

@@ 4.2 Adaptive clamped entropy control
where $\beta$ is the coefficient learning rate, and $\tilde{{\mathcal{H}}}_{\rm low},\tilde{{\mathcal{H}}}_{\rm high}$ are respectively the lower and upper limit of the (clamped) entropy. The algorithm will try to confine $\tilde{{\mathcal{H}}}$ within $\tilde{{\mathcal{H}}}_{\rm low},\tilde{{\mathcal{H}}}_{\rm high}]$ by increasing/decreasing $\lambda$ when $\tilde{{\mathcal{H}}}(\pi_{\theta})$ is lower/higher than the limits. The intuition is that when entropy is high, the coefficient should be tuned down to reduce the entropy induced bias and shift weights to reward maximization, which in turn consumes entropy. While when entropy level is too low, the coefficient can be tuned up to leverage the benefits of entropy regularization. For better training stability, the entropy coefficient is also boxed in the range $\lambda_{\rm low},\lambda_{\rm high}]$ so that large fluctuations of entropy do not lead to coefficient over-shoot. Empirically, we find that this scheme helps improve reasoning efficiency by avoiding entropy and response length explosion.
@@ 4.3 Algorithm
where ${\mathcal{L}}_{\rm PO}(\theta)$ is a policy optimization objective, e.g., the GRPO objective is used in our tests. At each global step, we set the sampling policy $\pi_{b}$ according to the choice of policy optimization objective ${\mathcal{L}}_{\rm PO}$. For example, in PPO-type algorithms, $\pi_{b}$ is set as the policy from last global step. Given $\pi_{b}$, a batch of queries $s_{0}\sim{\mathcal{D}}$ are sampled, and for each query, $\pi_{b}$ rolls out a batch of trajectories up to the maximum time step. With the batched samples, we can then optimize for the batch surrogate of ${\mathcal{L}}_{\rm AEnt}(\theta;\lambda)$ for several mini-epochs. At the end of each global step, the entropy coefficient is adjusted according to scheme 4.1. The whole process is summarized in Algorithm 1.
@@ 5.1 Training details
Evaluation. We evaluate models on the AIME 2024, MATH-Hard test split (Hendrycks et al., 2021), MATH-500 (Lightman et al., 2023), AMC23, MinervaMath (Lewkowycz et al., 2022) and OlympiadBench (He et al., 2024). We estimate the test score by averaging 4 tries per query on all benchmarks. The test-time generation temperature is $0.6$, top-p is $0.95$ and top-k is $20$.
Hyper-parameter settings. The tests are based on the verl framework (Sheng et al., 2025) [^1] The code will be released as soon as an internal mandatory pre-open-source review is completed.. For experiment setting (a), we use AdamW optimizer with a learning rate of $2\times 10^{-6}$. We set the max response length as $3072$. We use a batch size of $512$, and for each query we roll out $16$ responses. For AEnt, we use the GRPO objective as ${\mathcal{L}}_{\rm PO}$. We use a clamping percentage $p=0.33$, and set $\tilde{{\mathcal{H}}}_{\rm low}\!=\!0.15$ and $\tilde{{\mathcal{H}}}_{\rm high}\!=\!0.24$. We use an initial entropy coefficient $\lambda=0.002$, and start updating it from the third epoch with $\beta=0.002$. We clip the coefficient in between $0.0006$ and $0.009$. For EntReg method, we use the traditional entropy bonus with a fixed entropy coefficient of $0.002$. For experiment setting (b), we use a learning rate of $1\!\times\!10^{-6}$, a max response length of $7168$, a batch size of $256$ and for each query we roll out $8$ responses. We use $p=0.25$, $\tilde{{\mathcal{H}}}_{\rm low}\!=\!0.35$ and $\tilde{{\mathcal{H}}}_{\rm high}\!=\!0.62$, an initial $\lambda=3\!\times\!10^{-4}$, and start updating it from the second epoch with $\beta=10^{-4}$. We clip the coefficient in between $4\times 10^{-5}$ and $0.001$.
@@ 5.2 Performance analysis
An observation on the test score and the entropy trend. An interesting observation from Figures 4(a) is that after around $175$ steps (collapse time), the policy entropy of GRPO largely depletes and the entropy of EntReg starts to drastically fluctuate, while AEnt’s policy entropy is kept stable. Then one can observe from Figure 3(a) that the test score of GRPO and EntReg plateaus around the same step, while the score of AEnt continues to improve and surpasses the baselines past the collapse time. This observation is consistent with our intuition and theoretical analysis: after GRPO’s entropy collapse, its policy becomes concentrated on few paths and no new information can be gained in the sampling process, ultimately leading to the stagnancy of the learning process. This is predicted by Proposition (I) that the policy will become stationary after entropy depletion.
@@ 5.3 Ablation studies
In this section, we conduct ablation studies on our algorithm.
Adaptive coefficient stabilizes training. In Figure 5, we compare the performance of adaptive coefficient vs constant coefficient of the regularizer. The test performance is similar for the two methods in this particular experiment. However, adaptive coefficient leads to a significant advantage on reasoning efficiency by delivering more compact responses while not sacrificing accuracy. In the third plot of Figure 5, constant coefficient fails to stabilize policy entropy in the mid of training, which results in the entropy blow up. We observe a positive correlation between entropy and response length in this case, where a exploding entropy leads to repeated reasoning patterns that do not increase the test scores. On the other hand, the adaptive coefficient successfully prevents the entropy and response length from blowing up.
Analysis of the entropy clamping percentage $p$. In Figure 6, we compare the algorithmic performance under different choice of clamping percentage $p$. Intuitively, the percentage $p$ decides the size of the clamped space ${\mathcal{A}}(s)$, where a larger $p$ leads to more aggressive clamping and less tokens taken into account during entropy calculation. This would smooth the LLM policy on a more compact space, reducing the bias induced by entropy maximization while running the risk to leave out valuable tokens. In this sense, it is reasonable to try to maximize $p$ until the performance drops, which is also suggested by our reported results. Despite the fact the AEnt’s performance is affected by the choice of $p$, its advantage over the baselines is somewhat robust to the choice. It can be observed from Figure 6 that AEnt outperforms the baselines with different choices of $p$.
@@
5.4 Time complexity
We compare the time complexity of each algorithm under different base models and training datasets. The results are reported in Table 2. The experiments on Qwen-2.5-math-1.5b and MATH training dataset are conducted on 4xA100, and all other experiments are conducted on 8xA100. The hyper-paremter setting has been described in Section 5.1 where for fair comparison, we keep the batch sizes, the max response length and all computation speed related configures the same for all algorithms.
Qwen-math-1.5b+MATH R1-distilled-Qwen-1.5b+OpenR1
update per step to GRPO score to highest score update per step to GRPO score to highest score
GRPO 0.234 h 57 h 57 h 0.303 h 237 h 237 h
EntReg 0.253 h 49 h 52 h 0.316 h 215 h 238 h
AEnt 0.256 h 35 h 69 h 0.324 h 186 h 275 h
Table 2: Time complexity comparison under different settings. “Update per step” indicates the GPU hours of forward/backward process per step; “to GRPO/highest score” indicates the total GPU hours to reach the highest score achieved by GRPO/the algorithm itself. The first and second column respectively reports the results for setting (a) and (b) described in Section 5.1.
Overall, AEnt consumes slightly more compute per step due to 1) AEnt-trained model’s response length is moderately larger than those of the baselines, as indicated in Figure 4; 2) compared to GRPO, it requires the extra forward/backward process of the clamped entropy regularizer, and compared to EntReg, the additional clamping related computation is off-loaded to CPU in our implementation for memory save. However, AEnt reaches a common score threshold faster than the baselines as indicated by the “to GRPO score” results, indicating AEnt preserves the acceleration effect of entropy regularization. AEnt reaches the highest test score slower than the baselines since it prevents premature convergence, and converges to higher test scores which takes more training steps.
6 Conclusion
@@ Appendix A Appendix
A.1 Omitted notations in Section 2
Value functions. We first define some notations we omitted in Section 2. Given the definition of $V^{\pi_{\theta}})h(s)$, we can define the Q-function as $Q_{h}^{\pi_{\theta}}(s,a)=r(s,a)+V_{h+1}^{\pi_{\theta}}(s^{\prime})$ with $s^{\prime}=\mathcal{P}(s,a)$. By this definition, we can also equivalently write
$Q_{h}^{\pi_{\theta}}(s,a)=\mathbb{E}_{\pi_{\theta}}\Big[\sum_{t=h}^{H}r(s_{t},a_{t})|s_{h}=s,a_{h}=a\Big]$
We can also define the advantage function as $A_{h}^{\pi_{\theta}}(s,a)=Q_{h}^{\pi_{\theta}}(s,a)-V_{h}^{\pi_{\theta}}(s)$.
Entropy-regularized value functions. For the entropy-regularized setting, we can analogously define the entropy-regularized value functions as
$V_{h,\lambda}^{\pi_{\theta}}(s)\coloneqq\mathbb{E}_{\pi_{\theta}}\Big[\sum_{t=h}^{H-1}\big(r(s_{t},a_{t})-\lambda\log\pi_{\theta}(a_{t}|s_{t})\big)|s_{h}=s\Big]$
$Q_{h,\lambda}^{\pi_{\theta}}(s,a)\coloneqq r(s,a)+V_{h+1,\lambda}^{\pi_{\theta}}(s^{\prime})\text{ with }s^{\prime}=\mathcal{P}(s,a)$
Then the entropy-regularized advantage function is defined as $A_{h,\lambda}^{\pi_{\theta}}(s,a)=Q_{h,\lambda}^{\pi_{\theta}}(s,a)-\lambda\log\pi_{\theta}(a|s)-V_{h,\lambda}^{\pi_{\theta}}(s)$.
Note that we omit the time step subscript in the value functions when $h=0$, e.g., we write $V_{h}^{\pi_{\theta}}|_{h=0}$ as $V^{\pi_{\theta}}$ and similarly for all the value functions.
A.2 Preliminary lemmas
@@
A.3 Proof omitted in Section 3
A.3.1 Proof of Proposition 1
@@
Given $s_{0}$, by the assumption of our LLM tasks that $s_{t+1}=\mathcal{P}(s_{t},a_{t})$ is a concatenation of $s_{t},a_{t}$ for any $0\leq t\leq H-1$, we have $\mathbb{P}_{t}^{\pi_{\theta}}(s_{h}^{*})=0$ for any $t\neq h$. Using this fact in equation A.26
@@
$\|\nabla V^{\pi_{\theta}}({\mathcal{D}})\|$ $\geq C^{\pi_{\theta}}(s_{0})\Big(V^{\pi^{*}}(s_{0})-V^{\pi_{\theta}}(s_{0})\Big)$ (A.31)
where $C^{\pi_{\theta}}(s_{0})=\frac{1}{\sqrt{H}|{\mathcal{D}}|}\max_{(a_{0},\dots,a_{H-1})\in{\mathcal{A}}_{H}^{*}(s_{0})}\Pi_{t=0}^{H-1}\pi_{\theta}(a_{t}|s_{t})$ with ${\mathcal{A}}_{H}^{*}(s_{0})=\{(a_{0},a_{1},\dots,a_{H-1})\in{\mathcal{A}}^{H}~|~\exists\pi^{*}\in\arg\max_{\pi}V^{\pi}({\mathcal{D}}),\Pi_{t=0}^{H-1}\pi^{*}(a_{t}|s_{t})>0\}$. ∎
A.4 Proof of Proposition 2
@@
$V_{\lambda}^{\pi_{\lambda}^{*}}(s_{0})-V_{\lambda}^{\pi_{\theta}}(s_{0})\leq\frac{1}{2\lambda}\frac{1}{C_{\lambda}^{\pi_{\theta}}(s_{0})}\|\nabla V_{\lambda}^{\pi_{\theta}}({\mathcal{D}})\|^{2}$ (A.39)
@@
Define ${\mathcal{S}}(s_{0})\subseteq{\mathcal{S}}$ as the set of all possible states starting from $s_{0}$, i.e., ${\mathcal{S}}(s_{0})=\{s_{0}\}\cup\{s_{t}\in{\mathcal{S}}|t\in\{1,...,H-1\},a_{t-1}\in{\mathcal{A}},s_{t-1}\in{\mathcal{S}}(s_{0}),s_{t}=\mathcal{P}(s_{t-1},a_{t-1})\}$. Then we have
@@
where the second and the third inequalities follow from Cauchy-Schwartz inequality, and the last inequality follows from equation A.4. The constant $C_{d}=\frac{1}{\sqrt{|{\mathcal{S}}(s_{0})|}}$.
@@
$\geq C_{d}^{2}\sum_{s\in{\mathcal{S}}(s_{0})}\sum_{t=0}^{H-1}(\mathbb{P}_{t}^{\pi_{\theta}}(s)\min_{a}\pi_{\theta}(a|s))^{2}\Big\|Q_{t,\lambda}^{\pi_{\theta}}(s,\cdot)-\lambda\theta_{s,\cdot}-\frac{\sum_{a}Q_{t,\lambda}^{\pi_{\theta}}(s,a)-\lambda\theta_{s,a}}{|{\mathcal{A}}|}\Big\|_{\infty}^{2}$
$=C_{d}^{2}\lambda^{2}\sum_{s\in{\mathcal{S}}(s_{0})}\sum_{t=0}^{H-1}\frac{(\mathbb{P}_{t}^{\pi_{\theta}}(s)\min_{a}\pi_{\theta}(a|s))^{2}}{\mathbb{P}_{t}^{\pi_{\lambda}^{*}}(s|s_{0})}\mathbb{P}_{t}^{\pi_{\lambda}^{*}}(s|s_{0})\Big\|Q_{t,\lambda}^{\pi_{\theta}}(s,\cdot)/\lambda-\theta_{s,\cdot}$
@@
$C_{\lambda}^{\pi_{\theta}}(s_{0})$ $=C_{d}^{2}\min_{t,s\in{\mathcal{S}}(s_{0})}\frac{(\mathbb{P}_{t}^{\pi_{\theta}}(s)\min_{a}\pi_{\theta}(a|s))^{2}}{\mathbb{P}_{t}^{\pi_{\lambda}^{*}}(s|s_{0})}.$
@@
$V_{\lambda}^{\pi_{\lambda}^{*}}(s_{0})-V_{\lambda}^{\pi_{\theta}}(s_{0})\leq\frac{1}{2\lambda}\frac{1}{C_{\lambda}^{\pi_{\theta}}(s_{0})}\|\nabla V_{\lambda}^{\pi_{\theta}}({\mathcal{D}})\|^{2}$ (A.49)
which completes the proof. ∎
A.5 Toy verification in Figure 1
To verify the claim made after Proposition 2, we will need to vary the sparsity of optimal tokens and observe experimental results. However, it is generally difficult to control the sparsity of optimal responses for real queries. Thus the verification experiments reported in Figure 1 are conducted in a synthetic task, and we leave the results on non-synthetic tasks to Section 5.
Task setting. In the synthetic task, the total number of actions is $|{\mathcal{A}}|=10^{5}$, where the number of optimal action $n_{\rm opt}$ can be picked from $\{15,10,5,1\}$ and the number of suboptimal action is fixed at $500$. The reward for optimal action is $1$, for suboptimal action is $0.2$ and is 0 for all other actions. For simplicity, we set $H=1$. The policy is given by a tabular softmax with parameter $\theta\in\mathbb{R}^{10^{5}}$. To mimic a pre-trained initial policy, we initialize the policy parameter corresponding to the optimal actions and $500$ other random actions from $\mathcal{N}(1,1)$; while we initialize all other logits from $\mathcal{N}(0,1)$. Results in the figure are averaged over 20 independent runs.
Hyper-parameters. The learning rate is set as $0.02$, and the batch size is $64$. The hyper-parameters for each algorithm are found through a grid search. For $n_{\rm opt}=15,10,5,1$: the regularization coefficient for entropy regularization is set as $0.0005,0.0005,0.0005,0.0007$, and the coefficient for clamped entropy regularization is set as $0.0008$ uniformly. The clamping percentage $p$ is set as $0.98,0.98,0.985,0.997$.